Suppose that a One-way ANOVA is being performed to compare the means of 4 populations and that the sample sizes are 15, 17, 20, and 14. Determine the degrees of freedom for the F-statistic. (a) the degree of freedom of the numerator (b) the degree of freedom of the denominator
Answer #1

We have here,

k=4

n=15+17+20+14 =66

(a) the degree of freedom of the numerator=k-1=4-1=3

(b) the degree of freedom of the denominator =n-k=66-4=62